A 800mL container holds 2.366g of an unknown gas at 78.0 degrees Celsius and 103kPa. Identify the gas.
1
Expert's answer
2013-09-24T11:04:39-0400
Based on the ideal gas law the molecular mass of the gas was obtained in the following way:
PV = nRT; T = 351 K;
Then,
n = m/M = (PV)/(RT);
From here we determine
M = (mRT)/(PV) = (2.366*10-3*8.31*351)/(103*103*0.8*10-3) = 0.0837kg/mol = 83.7 g/mol
Therefore, the closest match is Krypton that has a molecular weight of 83.798 g/mol. Its boiling point is 119.93 K that is consistent with our
data as it's supposed to be a gas at 351K.